20. The following table presents data on imports and exports as a percentage of gross domestic product (GDP) for select countries in 2015, based on data from the World Bank. Exports as a Percentage of GDP Imports as a Percentage of GDP 21.2 82.7 30.3 18.8 38.9 15.5 Country Australia Belgium Chile China South Korea 45.9 United States 12.6 19.8 84.4 30.1 22.4 Based on data from the table, which countries had a trade deficit in 2015? (3 points)
Answer #1

"Trade Deficit" can be defined as a measure of trade by which imports exceeds exports.

In other words, when imports are higher as compared to exports, the country has trade deficit.

From the given data, imports as a percentage of GDP are higher for Australia, Chile and United States.

Australia = Exports (% of GDP) - Imports (% of GDP) = 19.8 - 21.2 = -1.4%

Chile = Exports (% of GDP) - Imports (% of GDP) = 30.1 - 30.3 = -0.2%

United States = Exports (% of GDP) - Imports (% of GDP) = 12.6 - 15.5 = -2.9%